# Early Classification of Time Series in Non-Stationary Cost Regimes

### v1

@@ I Motivation
I - In healthcare monitoring [11], for instance, the balance between misclassification and decision delay costs evolves as clinical protocols, hospital workload, or available treatments change. Early detection of patient deterioration enables timely intervention, yet premature alarms may lead to unnecessary treatments, inefficient resource allocation. Thus, the relative importance of accuracy versus earliness may drift over time, inducing a deployment-time objective that differs from the one assumed during training.
@@
  • Online adaptation of representative ECTS methods. We adapt several widely used ECTS approaches to the online setting, including an online extension of Calimera [6] based on a deep learning architecture. We also introduce (i) four new baselines—two based on multi-armed bandits and two adapting the probability-thresholding paradigm commonly used in classical ECTS—and (ii) propose lightweight online adaptations of Alert [25] and Economy [1], which were originally designed with online decision-making in mind but had not previously been evaluated in this context.
@@ II-C Online ECTS with deployment-time cost mismatch
To clearly distinguish these two temporal dimensions, we explicitly separate the time index of the time series from the deployment timeline. We denote by $\mathrm{C}^{(u)}_{m}$ and $\mathrm{C}^{(u)}_{d}$ the deployment-time misclassification and delay costs at deployment step $u\in\mathbb{N}^{+}$, where $u$ indexes successive processed incoming-time series. We define the corresponding loss as follows: $\mathcal{L}^{(u)}(\hat{y}(\mathbf{x}_{t}),y,t)=\mathrm{C}^{(u)}_{m}(\hat{y}(\mathbf{x}_{t})\mid y)+\mathrm{C}^{(u)}_{d}(t).$ In the general case, $\mathrm{C}^{(u)}_{m}$ and $\mathrm{C}^{(u)}_{d}$ are (i) unknown to the learner at deployment; (ii) may differ from the training-time costs ; (iii) and it can be non-stationary (drifting over time) and/or stochastic (varying across instances).
@@
  • Offline, a classifier is trained using full labeled time series and nominal costs.
@@ III-A Cost-Sensitive Learning
Learning algorithms that optimize decisions under user-defined cost functions have been extensively studied in cost-sensitive learning [10, 26, 23], typically by incorporating misclassification cost matrices into standard classifiers to minimize expected cost rather than error rate (e.g., via threshold adjustment, cost-sensitive losses, or tailored ensembles).
In an online setting, cost-sensitive learning has also been explored, e.g., [29] integrate fixed misclassification costs into streaming classifiers, but they do not address changes in the cost function over time. In contrast, the ECTS setting studied here optimizes an online time-dependent loss, combining misclassification and delay costs, under non-stationary deployment-time conditions.
@@ III-B Concept Drift in Data Streams
The scope of the present work is more specific. Rather than addressing changes in the data or label distribution, the focus is on cost drift: the underlying data-generating process is assumed stationary, while the decision costs governing the loss function may evolve. From this perspective, the proposed setting can be seen as an instance of learning under a fixed concept but evolving evaluation criteria.
@@ IV Bringing ECTS to Online Learning
The trigger function may be adaptive, learning from each processed time series and from the cost incurred by the corresponding decision. A key aspect of online adaptation is the update time, defined as the moment at which feedback becomes available to update the trigger. Depending on the algorithm and the information accessible after a decision, three update regimes can be distinguished: (i) Delayed updates correspond to settings where the learning signal is only available at time $T$, after the full time series has been observed. This occurs when the outcome associated with a decision can only be assessed once the entire process has unfolded. (ii) Instant updates refer to settings where the cost of the decision is immediately observable, allowing the trigger to be updated right after stopping. This regime naturally arises in RL approaches. Finally, (iii) No updates denotes methods in which the trigger is not adapted from feedback, but instead relies on explicit cost information specified at deployment time.
@@ IV-C No Updates (Cost-Agnostic Training)
Reproducibility: all the approaches presented in this paper are available as open-source code at: https://anonymous.4open.science/r/ects_cost_drift-BA28.
@@ V-D Evaluated Methods
Our experimental study compares the online ECTS triggering strategies introduced in Section IV, namely Proba-Threshold, Decay Proba-Threshold, HUCB1, SW-HUCB1, Deep-Calimera, Alert, and Economy-$\gamma$. In addition, two static baselines are included to provide reference points for evaluating the benefit of online adaptation: (i) no_adaptno\_adapt: a static Proba-Threshold model that does not adapt at deployment, consistently applying the threshold learned during training. (ii) silversilver: a Proba-Threshold model retrained from scratch at each deployment step $u$ using the true deployment-time cost configuration.
@@ VI-C Computation time
(decay)Proba-Threshold 0.076 2.546 2.622
@@ VII Conclusion
This paper addressed Early Classification of Time Series (ECTS) under deployment-time cost non-stationarity, a practically relevant setting that is largely overlooked in existing work. Beyond the specific ECTS setting, this work contributes to the broader problem of early and adaptive decision-making under evolving cost constraints.
We investigated two complementary sources of non-stationarity: deterministic drift in misclassification and delay costs, and stochastic deployment-time cost realizations. Focusing on separable ECTS approaches, we proposed online adaptations of representative triggering strategies that update only the trigger during deployment, while keeping the classifier fixed. This choice reflects realistic operational constraints and isolates the role of online cost adaptation.
@@
This work opens several directions for future research, including evaluation on real-world datasets, improving end-to-end ECTS methods to make it competitive, and jointly addressing cost non-stationarity, data distribution shifts and concept drifts. To facilitate further research, all experiments are fully reproducible and the implementation of the proposed methods is released as open source.
We believe that explicitly accounting for cost non-stationarity is a necessary step toward deploying reliable and adaptive neural decision systems in real-world, long-running applications.

### v2

@@ I Motivation
I - In healthcare monitoring [11], for instance, the balance between misclassification and decision delay costs evolves as clinical protocols, hospital workload, or available treatments change. Early detection of patient deterioration enables timely intervention; yet, premature alarms may lead to unnecessary treatments, inefficient resource allocation. Thus, the relative importance of accuracy versus earliness may drift over time, inducing a deployment-time objective that differs from the one assumed during training.
@@
  • Online adaptation of representative ECTS methods. We adapt several widely used ECTS approaches to the online setting, including an online extension of Calimera [6] based on a deep learning architecture. We also introduce (i) four new baselines: two based on multi-armed bandits and two adapting the probability-thresholding paradigm commonly used in classical ECTS and (ii) propose lightweight online adaptations of Alert [25] and Economy [1], which were originally designed with online decision-making in mind but had not previously been evaluated in this context.
@@ II-C Online ECTS with deployment-time cost mismatch
To clearly distinguish these two temporal dimensions, we explicitly separate the time index of the time series from the deployment timeline. We denote by $\mathrm{C}^{(u)}_{m}$ and $\mathrm{C}^{(u)}_{d}$ the deployment-time misclassification and delay costs at deployment step $u\in\mathbb{N}^{+}$, where $u$ indexes successive processed incoming-time series. We define the corresponding loss as follows: $\mathcal{L}^{(u)}(\hat{y}(\mathbf{x}_{t}),y,t)=\mathrm{C}^{(u)}_{m}(\hat{y}(\mathbf{x}_{t})\mid y)+\mathrm{C}^{(u)}_{d}(t).$ In the general case, $\mathrm{C}^{(u)}_{m}$ and $\mathrm{C}^{(u)}_{d}$ are (i) unknown to the learner at deployment; (ii) may differ from the training-time costs; (iii) and can be non-stationary (drifting over time) and/or stochastic (varying across instances).
@@
  • Offline, a classifier is trained using complete labeled time series and nominal costs.
@@ III-A Cost-Sensitive Learning
Learning algorithms that optimize decisions under user-defined cost functions have been extensively studied in cost-sensitive learning [10, 26, 23], typically by incorporating misclassification cost matrices into standard classifiers to minimize expected cost rather than the error rate (e.g., via threshold adjustment, cost-sensitive losses, or tailored ensembles).
In an online setting, cost-sensitive learning has also been explored, e.g., [29] integrate fixed misclassification costs into streaming classifiers; however, they do not address changes in the cost function over time. In contrast, the ECTS setting studied here optimizes an online time-dependent loss, combining misclassification and delay costs, under non-stationary deployment-time conditions.
@@ III-B Concept Drift in Data Streams
The scope of the present work is more specific. Rather than addressing changes in the data or label distribution, the focus is on cost drift: the underlying data-generating process is assumed to be stationary, while the decision costs governing the loss function may evolve. From this perspective, the proposed setting can be seen as an instance of learning under a fixed concept but evolving evaluation criteria.
@@ IV Bringing ECTS to Online Learning
The trigger function may be adaptive, learning from each processed time series and the cost incurred by the corresponding decision. A key aspect of online adaptation is the update time, defined as the moment at which feedback becomes available to update the trigger. Depending on the algorithm and the information accessible after a decision, three update regimes can be distinguished: (i) Delayed updates correspond to settings where the learning signal is only available at time $T$, after the full time series has been observed. This occurs when the outcome associated with a decision can only be assessed once the entire process has unfolded. (ii) Instant updates refer to settings where the cost of the decision is immediately observable, allowing the trigger to be updated right after stopping. This regime naturally arises in RL approaches. Finally, (iii) No updates denotes methods in which the trigger is not adapted from feedback, but instead relies on explicit cost information specified at deployment time.
@@ IV-C No Updates (Cost-Agnostic Training)
Reproducibility: all the approaches presented in this paper are available as open-source code at Github.
@@ V-D Evaluated Methods
Our experimental study compares the online ECTS triggering strategies introduced in Section IV, namely Proba-Threshold, decay-Proba-Threshold, HUCB1, SW-HUCB1, Deep-Calimera, Alert, and Economy-$\gamma$. In addition, two static baselines are included to provide reference points for evaluating the benefit of online adaptation: (i) no_adaptno\_adapt: a static Proba-Threshold model that does not adapt at deployment, consistently applying the threshold learned during training. (ii) silversilver: a Proba-Threshold model retrained from scratch at each deployment step $u$ using the true deployment-time cost configuration.
@@ VI-C Computation time
(decay-)Proba-Threshold 0.076 2.546 2.622
@@ VII Conclusion
This paper addresses ECTS under deployment-time cost non-stationarity, a practically relevant setting that is largely overlooked in existing work. Beyond the specific ECTS setting, this work contributes to the broader problem of early and adaptive decision-making under evolving cost constraints.
We investigated two complementary sources of non-stationarity: deterministic drift in misclassification and delay costs, and stochastic deployment-time cost realizations. Focusing on separable ECTS approaches, we proposed online adaptations of representative triggering strategies that update only the trigger during deployment.
@@
This work opens several directions for future research, including evaluation on real-world datasets, improving end-to-end ECTS methods to make it competitive, and jointly addressing cost non-stationarity, data distribution shifts and concept drifts. To facilitate further research, all experiments are fully reproducible and are released as open source. We believe that explicitly accounting for cost non-stationarity is a necessary step toward deploying reliable and adaptive neural decision systems in real-world, long-running applications.